A particle is moving along a straight line such that its acceleration can be given as a=3-4t, where t is in second. At t=0, its velocity is v=2m/s. Distance travelled by particle in 3 s is

Grade:11

1 Answers

Aman
6 years ago
Using a=dv/dt v=3t+2-2t^2 taking lt of v as 0-2 & that of as 0 -t. Solving v=0 t comes two. Using v=dx/dt x=-2/3 tcube+ 3/2 tsq +2t . x in t=2 is 14/3 and x in 3 is 3/2. As at t=2 v change direction total distance would be 2× x in t=2 - x in t=3= 47/6